Find a value for n to ensure that the absolute error in approximating the integral
∫_0^2▒〖x cos⁡〖x dx〗 〗by the midpoint rule will be less than 10^-3
1
Expert's answer
2020-10-21T14:26:53-0400

Absolute error in midpoint approximation is calculated as "\frac{(b-a)^3k}{24n^2}" , where k is a max value of |f''(x)| on [a,b].


In this care a=0, b=2, f(x)=xcos(x), f'(x)=cos(x)-xsin(x), f''(x)=-sin(x)-sin(x)-xcos(x)=-2sin(x)-xcos(x), f'''(x)=-2cos(x)-cos(x)+xsin(x)=-3cos(x)+xsin(x), f''''(x)=3sin(x)+sin(x)-xcos(x)=4sin(x)+xcos(x).


-3cos(x)+xsin(x)=0


xsin(x)=3cos(x)


x=3cot(x). Using a numerical method, we get that x=1.1925 on [0,2].


|f''(0)|=|-2sin(0)-0|=0.


|f''(1.1925)|=|-2sin(1.1925)-1.1925cos(1.1925)|=2.299


|f''(2)||=|-2sin(2)-2cos(2)|=0.9863.


Therefore, a max value of |f''(x)| on [0,2] is 2.299


So, if absolute error is less than 0.001 then:


"\frac{2^3\cdot2.299}{24n^2}\leq0.001

\newline

\newline


24n^2\cdot0.001\geq2^3\cdot2.299

\newline

n^2\geq\frac{2^3\cdot2.299}{24\cdot0.001}

\newline 

n^2\geq 766.33

\newline

n\geq \sqrt{766.33}

\newline

n\geq 27.6827"


So n has to be at least 28.


Answer: n=28